Find the cylindrical coordinates of the points where the Cartesian coordinates are

i.(6,6,8)

ii.(√2,1,1)


1
Expert's answer
2021-02-04T08:09:30-0500

Solution (i):

Given, "x=6,\ y=6, z=8"

In rectangular coordinates,

"x=r\cos\theta,\ y=r\sin\theta,\ z=z"

So, "6=r\cos\theta,\ 6=r\sin\theta,\ z=8"

Also, "r^2=x^2+y^2=6^2+6^2=36+36=72"

"\Rightarrow r=\sqrt{72}=6\sqrt2" (Taking positive value only)

And, "\tan\theta=\dfrac yx=\dfrac 66=1"

"\Rightarrow \theta=\tan^{-1}1=\dfrac{\pi}4"

So, cylindrical coordinates are "(r,\theta,z)=(6\sqrt2,\dfrac {\pi}4,8)"

Solution (ii):

Given, "x=\sqrt2,\ y=1, z=1"

In rectangular coordinates,

"x=r\cos\theta,\ y=r\sin\theta,\ z=z"

So, "\sqrt2=r\cos\theta,\ 1=r\sin\theta,\ z=1"

Also, "r^2=x^2+y^2=(\sqrt2)^2+1^2=2+1=3"

"\Rightarrow r=\sqrt3" (Taking positive value only)

And, "\tan\theta=\dfrac yx=\dfrac 1{\sqrt2}"

"\Rightarrow \theta=\tan^{-1}(\dfrac 1{\sqrt2})=0.6154" rad

So, cylindrical coordinates are "(r,\theta,z)=(\sqrt3,0.6154,1)"